(a) Is every total ordering a lattice? Why or why not?
1
Expert's answer
2020-12-15T02:27:25-0500

A partially ordered set "(L, \u2264)" is called a lattice if each two-element subset "\{a, b\} ⊆ L" has supremum and infimum in "L", denoted by "a \u2228 b" and "a \u2227 b", respectively.


If "(L, \u2264)" is a total ordering, then by definition, "{\displaystyle a\leq b}" or "{\displaystyle b\leq a}" for all "a,b\in L". If "{\displaystyle a\leq b}", then "a ∨ b=b\in L" and "a ∧ b=a\in L". If "{\displaystyle b\leq a}", then "a ∨ b=a\in L" and "a ∧ b=b\in L".


Therefore, every total ordering is a lattice.